OpenStudy (anonymous):

Find the area that lies between the two circles r = 10sin(x) r = 10cos(x) This is a calc 2 polar question

OpenStudy (tkhunny):

Well, then. I guess you'll need an integral in Polar Coordinates. What's your plan?

OpenStudy (anonymous):

If I recall correctly, aren't you supposed to use something like: \[\large \int_\alpha ^\beta \frac{1}{2}[r(\theta)]^2d\theta \]

OpenStudy (anonymous):

I would highly recommend drawing a picture first.

OpenStudy (anonymous):

Sorry it took so long for me to respond my internet went down. My picture is as follows.

OpenStudy (anonymous):

as for my integral I was looking to use\[10\int\limits_{0}^{\pi/4}\sin(\theta)+10\int\limits_{\pi/4}^{\pi/2}\cos(\theta)\]

OpenStudy (anonymous):

I know the area that should be used for polar is \[\int\limits_{a}^{b} \pi (f(r))^2 d \theta\]

OpenStudy (anonymous):

I wasn't sure if I should use the area equation for each integral

OpenStudy (tkhunny):

Why would you use anything else?